OpenStudy (anonymous):

solve 1,296^2 = 216^n+10

OpenStudy (anonymous):

@waterineyes i dont think its a quadratic equation .. \[1296^{2} - 10 = 216^{n}\] taking log on both side \[\ln (1296^{2}-10)=\ln (216^{n})\]

OpenStudy (anonymous):

Oh sorry..

OpenStudy (anonymous):

\[1296^2 = 216^n + 10\] \[1679616 - 10 = 216^n\] Taking Log both the sides, \[Log(1679606) = Log(216)^n\] \[6.225 = n \times Log(216)\] \[6.225 = n \times 2.334\] \[n = \frac{6.225}{2.334}\] \[\huge \color{Cyan}{n = 2.67}\]